An oil drop of mass 1.0mg and having a charge of 2.0 x 10^-6 is suspended in a box. What is the magnitude of the electric field in the box?


1
Expert's answer
2021-04-23T11:43:17-0400

Mass of oil drop, "m=1\space mg=1\times10^{-6}\space kg"

Charge, "q=2\times10^{-6}\space C"

"F=qE=mg"

"E=\dfrac{mg}{q}"

"E=\dfrac{10^{-6}\times 9.8}{2\times10^{-6}}=4.9\space N/C(upwards)"